# On-Robot Reinforcement Learning with Goal-Contrastive Rewards

### v1

@@ Abstract
Reinforcement Learning (RL) has the potential to enable robots to learn from their own actions in the real world. Unfortunately, RL can be prohibitively expensive, in terms of on-robot runtime, due to inefficient exploration when learning from a sparse reward signal. Designing dense reward functions is labour-intensive and requires domain expertise. In our work, we propose Goal-Contrastive Rewards (GCR), a dense reward function learning method that can be trained on passive video demonstrations. By using videos without actions, our method is easier to scale, as we can use arbitrary videos. GCR combines two loss functions, an implicit value loss function that models how the reward increases when traversing a successful trajectory, and a goal-contrastive loss that discriminates between successful and failed trajectories. We perform experiments in simulated manipulation environments across RoboMimic and MimicGen tasks, as well as in the real world using a Franka arm and a Spot quadruped. We find that GCR leads to a more-sample efficient RL, enabling model-free RL to solve about twice as many tasks as our baseline reward learning methods. We also demonstrate positive cross-embodiment transfer from videos of people and of other robots performing a task. Appendix: https://tinyurl.com/gcr-appendix-2.
@@ VI EXPERIMENTS
Figure 5: (Left) Cumulative returns of DrQ trained with different reward functions (x axis). We normalize the cumulative returns by the highest achieved value across all methods and runs, since the absolute returns are not comparable across tasks. We report four random seeds across three simulated tasks, SERL Pick cube, RoboMimic Lift and Can, and MimicGen Stack D0. (Right) We also add results for real-world Franka Lift, Kettle and Spot drawer opening tasks (Figure 6). We run these tasks with only GCR and Sparse rewards.
@@
Baselines and ablations: We compare our method to previous Inverse Reinforcement Learning methods. Specifically, Rank2Reward (R2R) [38] and a version of Generative Adversarial Imitation Learning (GAIL) [35], Variational Inverse Control with Events (VICE) [36] and Adversarial Inverse Reinforcement Learning (AIRL) [36] adapted to passive videos. We further compare GCR to only using sparse rewards (Sparse) and to using sparse rewards with RLPD [4] (Sparse + RLPD). RLPD requires demonstrations with actions. We consider five ablations of our method. VIP Frozen: a reward function derived from VIP fine-tuned on passive demonstrations and frozen during RL training; VIP Online: VIP fine-tuned on both passive demonstrations and online data from the RL agent; GCR (abs) uses the absolute reward (Section V-B, $\alpha=0,\beta=1$) instead of a delta reward; GCR (IC) uses an InfoNCE contrastive loss instead of the Simple Contrastive (GCR (SC)) loss. We use an unfrozen ResNet-50 LIV backbone [15] pre-trained on EPIC-KITCHENS [19] in all our experiments unless stated otherwise. Further details are in our appendix [^1] Appendix: https://tinyurl.com/gcr-appendix-2.

### v2

@@ Abstract
Reinforcement Learning (RL) has the potential to enable robots to learn from their own actions in the real world. Unfortunately, RL can be prohibitively expensive, in terms of on-robot runtime, due to inefficient exploration when learning from a sparse reward signal. Designing dense reward functions is labour-intensive and requires domain expertise. In our work, we propose Goal-Contrastive Rewards (GCR), a dense reward function learning method that can be trained on passive video demonstrations. By using videos without actions, our method is easier to scale, as we can use arbitrary videos. GCR combines two loss functions, an implicit value loss function that models how the reward increases when traversing a successful trajectory, and a goal-contrastive loss that discriminates between successful and failed trajectories. We perform experiments in simulated manipulation environments across RoboMimic and MimicGen tasks, as well as in the real world using a Franka arm and a Spot quadruped. We find that GCR leads to a more-sample efficient RL, enabling model-free RL to solve about twice as many tasks as our baseline reward learning methods. We also demonstrate positive cross-embodiment transfer from videos of people and of other robots performing a task. Website: https://gcr-robot.github.io/.
@@ VI EXPERIMENTS
Figure 5: (Left) Cumulative returns of DrQ trained with different reward functions (x axis). We normalize the cumulative returns by the highest achieved value across all methods and runs. We report four random seeds across three simulated tasks, SERL Pick cube (20 passive demos), RoboMimic Lift (20) and Can (20), and MimicGen Stack D0 (100). (Right) We also add results for real-world Franka Lift, Kettle and Spot drawer opening tasks (Figure 6). We run these tasks with only GCR and Sparse rewards.
@@
Baselines and ablations: We compare our method to previous Inverse Reinforcement Learning methods. Specifically, Rank2Reward (R2R) [38] and a version of Generative Adversarial Imitation Learning (GAIL) [35], Variational Inverse Control with Events (VICE) [36] and Adversarial Inverse Reinforcement Learning (AIRL) [36] adapted to passive videos. We further compare GCR to only using sparse rewards (Sparse) and to using sparse rewards with RLPD [4] (Sparse + RLPD). RLPD requires demonstrations with actions. We consider five ablations of our method. VIP Frozen: a reward function derived from VIP fine-tuned on passive demonstrations and frozen during RL training; VIP Online: VIP fine-tuned on both passive demonstrations and online data from the RL agent; GCR (abs) uses the absolute reward (Section V-B, $\alpha=0,\beta=1$) instead of a delta reward; GCR (IC) uses an InfoNCE contrastive loss instead of the Simple Contrastive (GCR (SC)) loss. We use an unfrozen ResNet-50 LIV backbone [15] pre-trained on EPIC-KITCHENS [19] in all our experiments unless stated otherwise. Further details are in the Appendix APPENDIX.
@@ ACKNOWLEDGMENT
APPENDIX
VII-A Simulated Environment Details
We adapt SERL Pick cube [25], RoboMimic Lift, Can, and Square [26], and MimiGen Stack D0 and Coffee D0 [27] environments to a sparse reward setting, where the agent receives a reward of 0 until reaching a goal condition, which has a reward of 1. The environments have a fixed episode length: 100 in Pick cube and 400 in RoboMimic Lift, Can, and Square, and 800 in MimicGen Stack D0 and Coffee D0. We use a pair of shoulder and wrist RGB cameras as inputs to the RL agent; the reward predictors only use the shoulder camera. The action space in all environments consists of the delta motions of the end-effector (EE) at approximately 5 Hz as well as a gripper open/closed fraction command. All environments are instantiated with a Franka robot arm. In Pick cube, the EE rotation is fixed and in Lift, the EE rotation is unconstrained. We simplify the RoboMimic Can and Square environments by fixing the EE rotation in Can and limiting EE rotation to yaw only in Square. We use the provided demonstrations in Pick cube, Stack D0 and Coffee D0 and collect a new dataset of teleoperated demonstrations in Lift, Can, and Square using a 3Dconnexion SpaceMouse to teleoperate the robot. We also collect demonstrations in Lift and Square with a Kuka robot instead of a Franka.
VII-B Real Environment Details
Figure 9: Our Boston Dynamics Spot setup used in the drawer opening experiment.
Figure 10: Our Franka tabletop setup used in the plushie picking and kettle opening experiments.
We perform real-world experiments with a table-top Franka arm (Figure 10) and a Boston Dynamics Spot with a mounted arm (Figure 9). The Franka task is to pick up a plushie and open a kettle by pushing a button. The Spot task is to open a drawer. We use a front-facing and a wrist-mounted camera in the Franka setup and a fixed EE rotation. The Spot setup involves two cameras mounted on the front of the Spot and the wrist of the Spot arm. In both cases, the front camera is a RealSense D455 and the wrist camera D435. The Spot uses a Fin-ray style soft gripper; the EE rotation is fixed to face the drawer. The Spot base moves in a semicircle around the drawer after each episode. The Franka and Spot tasks have a fixed episode length of 100 and 150 respectively. The Spot experiment has a scripted policy for opening a drawer to collect demonstrations and for closing the drawer for environment resets. The Spot automatically sits down and rolls over when its battery is at 10% charge, after about an hour of running. We then swap the battery and resume the experiment. We use a SpaceMouse teleoperation setup to collect demonstrations with a Franka and reset its environment manually.
We also implement a goal classifier for each task, which provides extrinsic rewards to the reinforcement learning agent. For Spot drawer, we use an external camera to observe the drawer from the side. Then, we use Grounding DINO [63] to detect the drawer and Segment Anything [64] to predict its segmentation mask. We assign a reward of 1 if the mask has increased by a certain size throughout the episode, which happens only when the drawer is open. For Franka plushie, we similarly detect and segment a plushie. Then, we threshold the height of the segmentation mask to predict success. For Franka kettle, we ask GPT-4o vision to answer a yes/no question for each frame with the following prompt: ”Is the kettle open? Think it step by step but limit your answer to ONLY a word ’yes’ or ’no’. There might be a robot gripper inside the view but you should focus on the kettle. If you are not sure, answer ’no.’, no false positive is allowed. Only yes or no should appear in the answer. I will tip you \$500 for a correct answer.”
VII-C Implementation Details
We re-implement LIV [15] and a version of VIP [14] that uses improvements from LIV (most importantly, cosine distance instead of $L_{2}$ distance on embeddings). We initialize LIV with a ResNet-50 CLIP model and further pre-train it on the EPIC-KITCHENS dataset [19]. We then use the VIP loss (LIV without language) to fine-tune the model on each task. All methods use the LIV backbone pre-trained on EPIC-KITCHENS and fine-tuned during online learning. We further run Rank2Reward with the LIV and R3M backbone, both frozen and unfrozen, and report the best result. We use a SERL implementation of DrQ and RLPD [69] RL agents for all experiments. The agents are trained with a standard replay buffer and the Adam optimizer. We use the default hyper-parameters of DrQ and RLPD from SERL.
We train GCR and all baselines with a batch size of 32, a learning rate of $10^{-6}$, a weight decay of $10^{-3}$, and a negative buffer size of 20k steps. We use the GCR loss with $\omega_{1}=\omega_{2}=1$. We further multiply both the intrinsic and extrinsic rewards by 10, as this appears to help the RL agent learn faster from the first few times it experiences a success. For simulated experiments, we use two NVIDIA V100 GPUs for each random seed. The first GPU runs the RL actor, RL learner and intrinsic reward predictor. The second GPU runs the intrinsic reward learner. We use two NVIDIA A6000 GPUs in our real-world experiments. In these experiments, we run the RL actor and RL learner on the first GPU and the rest – extrinsic reward predictor, intrinsic reward predictor and intrinsic reward learner – on the second GPU.
We list the hyper-parameters of GCR, baselines and the deep RL agent in Table II.
TABLE II: GCR and baseline hyper-parameters used in our simulation experiments.
Parameter Setting
GCR
Batch size 32
Learning rate $10^{-6}$
Weight decay $10^{-3}$
Negative buffer size 20000
Negative alpha 0.95
$\omega_{1}$ 1
$\omega_{2}$ 1
VIP / LIV
Number of negative examples 3
Alpha 0.95
Metric cosine
Learning rate $10^{-6}$
Weight decay $10^{-3}$
Vision weight 1
CLIP weight 0
Language weight 0
L1 regularization weight 0
L2 regularization weight 0
Discount 0.98
Rank2Reward (+ GAIL, VICE, AIRL)
Batch size 256 (R3M Frozen)
128 (LIV Frozen)
32 (R3M Unfrozen)
16 (LIV Unfrozen)
Learning rate $10^{-4}$
Use mixup loss True
Use state noise True
DrQ / RLPD
Buffer size 200000
Batch size 256
UTD ratio 4
Random start steps 1000
Critic ensemble size 10
Critic subsample size 2
Discount 0.96
Initial temperature $10^{-2}$
Learning rate $3*10^{-4}$
Soft target update rate $5*10^{-3}$
Actor / learner steps $\sim$ 0.75